STRUCTURE DETERMINATION OF RESORCINOL ROTAMERS BY HIGH RESOLUTION UV SPEC-TROSCOPY

GRZEGORZ MYSZKIEWICZ, \underline{W. LEO MEERTS}, Molecular- and Biophysics Group, Institute for Molecules and Materials, Radboud University Nijmegen, NL-6500 GL Nijmegen, The Netherlands; CHRISTIAN RATZER, and MICHAEL SCHMITT, Heinrich-Heine-Universität, Institut für Physikalische Chemie, 40225 Düsseldorf, Germany.

Several rotationally resolved S_{1} ← S_{0} electronic origins of deuterated resorcinol rotamers cooled in a molecular beam were recorded. An automated assignment of the observed spectra was performed using a genetic algorithm approach with an asymmetric rotor Hamiltonian. The structure of resorcinol A and resorcinol B was derived from the rotational constants of several deuterated species for both electronic states. The lifetimes of different resorcinol isotopomers in the S_{1} state are also reported. Like in phenol these lifetimes mainly depend on the position of deuteration. A nearly perfect additivity of the zero-point energies after successive deuterations in resorcinol rotamers was discovered and subsequently used in the full assignment of the previously reported low resolution spectra of deuterated resorcinol A^{a}. The analogous spectrum was also predicted for the resorcinol B rotamer.

aM. Gerhards, C. Unterberg, and S. Schumm, J. Chem. Phys., 1999, 111, 7966–7975.